# Neural variational inference for cutting feedback during uncertainty propagation

### v1

@@ Appendix A Derivations for Table 1: Spikiness, conditional perturbation rate, and quantile growth
$\begin{split}L_{S}(M)=\sup_{\eta\in K_{\eta}}\ \sup_{\begin{subarray}{c}\theta\neq\theta^{\prime}\\ |\theta\,|\,,|\theta^{\prime}|\leq M\end{subarray}}\frac{|\log p(\theta\mid\eta)-\log p(\theta^{\prime}\mid\eta)|}{|\theta-\theta^{\prime}|},\\ L_{CP}(M)=\sup_{|\theta\,|\,\leq M}\ \sup_{\eta\neq\eta^{\prime}}\frac{|\log p(\theta\mid\eta)-\log p(\theta\mid\eta^{\prime})|}{\|\eta-\eta^{\prime}\|_{1}}.\end{split}$
For $G(M)$ we evaluate upper envelopes of $Q_{\eta}$ at the Gaussian tail levels $\varepsilon_{N}:=\bar{\Phi}(M)=1-\Phi(M)\sim\phi(M)/M$. On $K_{\eta}$ (compact), continuous parameter maps are bounded; for $L_{CP}(M)$ we use that $\log p(\theta\mid\eta)$ is smooth in its parameters on compact sets and apply the chain rule with the (finite) Lipschitz moduli of the parameter maps induced by continuity on $K_{\eta}$. [^2] For standard parametric families, the derivatives of $\log p$ with respect to the parameters are bounded on compacts; combined with bounded variation of the parameter maps on $K_{\eta}$, this gives the displayed $L_{CP}(M)$ orders.
@@
$Q_{\eta}(u)=\mu(\eta)+\sigma(\eta)\Phi^{-1}(u)\ \Rightarrow\ T(\eta,z)=\mu(\eta)+\sigma(\eta)z,\quad G(M)=\sup_{\eta}\,|\,\mu(\eta)|+\sup_{\eta}\sigma(\eta)\,N=O(M).$
$\partial_{\theta}\log p=-(\theta-\mu)/\sigma^{2}\ \Rightarrow\ L_{S}(M)=O\!\Big(\frac{M}{\sigma^{2}_{\min}}\Big)=O(M).$
@@
For $Z\sim\mathcal{N}(0,1)$, by Mill’s ratio, the upper Gaussian tail satisfies $\varepsilon_{N}=\Phi(-N)\sim\frac{\phi(M)}{M}.$
@@
So, $Q_{\eta}(\Phi(-z))=O(1/\Phi(z))=z\exp(z^{2}/2)$ as $z\to\infty$, giving the bound
$G(M)=O\!\left(\frac{M}{\phi(M)}\right)=O\!\left(Me^{M^{2}/2}\right).$
The log–density derivatives are
$\partial_{\theta}\log p=-\frac{2(\theta-\mu)}{\gamma^{2}+(\theta-\mu)^{2}},\qquad\partial_{\mu}\log p=\frac{2(\theta-\mu)}{\gamma^{2}+(\theta-\mu)^{2}},\qquad\partial_{\gamma}\log p=-\frac{1}{\gamma}+\frac{2(\theta-\mu)^{2}}{\gamma\big(\gamma^{2}+(\theta-\mu)^{2}\big)}.$
These satisfy the uniform bounds
@@
Student-$t_{\nu(\eta)}$ with location $\mu(\eta)$, scale $s(\eta)$. Tail $\bar{F}_{\eta}(x)\sim c_{\nu(\eta)}\,x^{-\nu(\eta)}$ yields
$Q_{\eta}(1-\varepsilon)\sim(c_{\nu(\eta)}/\varepsilon)^{1/\nu(\eta)}\ \Rightarrow\ G(M)=O\!\Big(\Big(\frac{M}{\phi(M)}\Big)^{1/\nu_{\min}}\Big)=O\!\big((Me^{M^{2}/2})^{1/\nu_{\min}}\big),$
@@
$\partial_{\theta}\log p=-\frac{(\nu+1)(\theta-\mu)}{\nu s^{2}+(\theta-\mu)^{2}}\ \Rightarrow\ L_{S}(M)=O(1).$
Derivatives w.r.t. $(\mu,s)$ are bounded on compacts as above; the $\nu$–term contains $\tfrac{1}{2}\log\!\big(1+(\theta-\mu)^{2}/(\nu s^{2})\big)$, hence on [M,M][-M,M] gives $O(\log M)$:
@@
With normal base $Z\sim\mathcal{N}(0,1)$, let $\varepsilon_{N}=\Phi(-N)\sim\phi(M)/M$. Then
@@
Log–density derivatives. For $\theta\neq\mu$,
@@
Uniform bounds on the $\theta$–truncated set $\{|\theta-\mu|\leq M\}$:
@@
Symmetrized mollified log–normal.
@@
Let $\mu(\eta),\sigma(\eta)$ be continuous on compact $K_{\eta}$ with $0<\sigma_{\min}\leq\sigma(\eta)\leq\sigma_{\max}<\infty$. Define the symmetrized log–normal density
$f_{\mathrm{SLN}}(\theta\mid\eta)=\tfrac{1}{2}f_{\mathrm{LN}}(|\theta\,|\,\mid\eta),\qquad f_{\mathrm{LN}}(x\mid\eta)=\frac{1}{x\sigma(\eta)\sqrt{2\pi}}\exp\!\Big(-\frac{(\log x-\mu(\eta))^{2}}{2\sigma^{2}(\eta)}\Big),\quad x>0.$
Let $K_{h}$ be a symmetric $C^{\infty}$ mollifier with compact support [h,h][-h,h], $\int K_{h}=1$, and $K_{h}(u)\geq\kappa_{0}>0$ for $|u|\leq h/2$. Define the smoothed symmetric log–normal
@@
Since $f_{\mathrm{LN}}(\cdot\mid\eta)$ is decreasing for $x$ beyond its mode, the worst case over $|\theta\,|\,\leq M$ is attained at $|\theta\,|\,=N$. Thus, for all $\eta\in K_{\eta}$,
$\inf_{|\theta\,|\,\leq M}p_{h}(\theta\mid\eta)\ \geq\ \frac{\kappa_{0}}{2}\!\int_{N-h/2}^{M+h/2}f_{\mathrm{LN}}(x\mid\eta)\,dx\ \geq\ \frac{\kappa_{0}h}{4}\,\inf_{\eta\in K_{\eta}}\ \inf_{x\in[N-h/2,\,M+h/2]}f_{\mathrm{LN}}(x\mid\eta).$
@@
$L_{S}(M)$ $=O\!\Big((M+1)\exp\!\Big(\tfrac{(\log(M+h/2)-\mu_{\max})^{2}}{2\sigma_{\min}^{2}}\Big)\Big)=O(Me^{O(\log M)^{2}}),$
$L_{CP}(M)$ $=O\!\Big((\log M)^{2}(M+1)\exp\!\Big(\tfrac{(\log(M+h/2)-\mu_{\max})^{2}}{2\sigma_{\min}^{2}}\Big)\Big)=O((\log M)^{2}Ne^{O(\log M)^{2}}).$
Median and derivative envelope. Symmetry gives $M_{0}=0$. Compact support of $K_{h}$ gives two–sided tail comparison for large $x$,
$c_{1}\,f_{\mathrm{SLN}}(x-h\mid\eta)\ \leq\ p_{h}(x\mid\eta)\ \leq\ c_{2}\,f_{\mathrm{SLN}}(x+h\mid\eta),$
with $c_{1},c_{2}>0$ depending only on $K_{h}$. Convolution and Fubini give the exact identity
$\bar{F}_{h,\eta}(x)=\int_{\mathbb{R}}K_{h}(t)\,\bar{F}_{\eta}(x-t)\,dt$
where $\overline{F}_{\eta}$ is the survival function for $f_{SLN}$ and $\overline{F}_{h,\eta}$ is that for $p_{h}$. Since $\bar{F}_{\eta}$ is nonincreasing and $t\in[-h,h]$ on the support of $K_{h}$, it follows that
$\bar{F}_{\eta}(x+h)\ \leq\ \bar{F}_{h,\eta}(x)\ \leq\ \bar{F}_{\eta}(x-h)\qquad\text{for all }x\in\mathbb{R}.$
Write the generalized quantile functions
$Q_{\eta}(u):=\inf\{x:\ F_{\eta}(x)\geq u\},\qquad Q_{h,\eta}(u):=\inf\{x:\ F_{h,\eta}(x)\geq u\},\qquad u\in(0,1).$
Equivalently, if $S_{\eta}:=\bar{F}_{\eta}$ and $S^{-1}_{\eta}(y):=\inf\{x:\ S_{\eta}(x)\leq y\}$, then $Q_{\eta}(u)=S^{-1}_{\eta}(1-u)$.
Fix $u\in(0,1)$ and set $x_{u}:=Q_{h,\eta}(u)$. Continuity of $F_{h,\eta}$ implies $\bar{F}_{h,\eta}(x_{u})=1-u$. The survival bounds at $x=x_{u}$ give
[…]
By monotonicity of $S_{\eta}$ and of its generalized inverse,
$x_{u}-h\ \leq\ S^{-1}_{\eta}(1-u)\ \leq\ x_{u}+h.$
Since $S^{-1}_{\eta}(1-u)=Q_{\eta}(u)$, we conclude
$\big|\,Q_{h,\eta}(u)-Q_{\eta}(u)\,\big|\ \leq\ h\qquad\text{for all }u\in(0,1).$
Setting $u=\Phi(z)$ yields
$\big|\,T_{h}(\eta,z)-T(\eta,z)\,\big|\ \leq\ h\qquad\text{for all }z\in\mathbb{R},$
with a constant $h$ that depends only on the support radius of $K_{h}$ and is independent of $\eta$.
For $x>0$,
$f_{\mathrm{LN}}(x;m,\sigma)=\frac{1}{x\,\sigma\sqrt{2\pi}}\exp\!\Big(-\frac{(\log x-m)^{2}}{2\sigma^{2}}\Big).$
Write $x_{z}:=T(\eta,z)=\exp\!\big(m(\eta)+\sigma(\eta)z+O(1)\big)$ as $|z|\to\infty$. Using the tail comparison at $x=T_{h}(\eta,z)=x_{z}+O(1)$,
$\log p_{h}\bigl(T_{h}(\eta,z)\mid\eta\bigr)=-\log f_{\mathrm{SLN}}\bigl(x_{z}+O(1)\mid\eta\bigr)+O(1).$
Since $f_{\mathrm{SLN}}(x\mid\eta)=\tfrac{1}{2}f_{\mathrm{LN}}(x;m(\eta),\sigma(\eta))$ for $x>0$,
$\log f_{\mathrm{SLN}}(x_{z}+O(1)\mid\eta)=\frac{\big(\log x_{z}-m(\eta)+O(x_{z}^{-1})\big)^{2}}{2\sigma(\eta)^{2}}+\log x_{z}+\log\sigma(\eta)+O(1).$
With $\log x_{z}=m(\eta)+\sigma(\eta)z+O(1)$ this becomes
$\log f_{\mathrm{SLN}}(x_{z}+O(1)\mid\eta)=\frac{z^{2}}{2}+\frac{O(1)}{\sigma(\eta)}\,|z|+\log x_{z}+O(1)=\frac{z^{2}}{2}+\sigma(\eta)\,z+O(1),$
uniformly over $\eta\in K_{\eta}$. Therefore
$\log p_{h}\bigl(T_{h}(\eta,z)\mid\eta\bigr)=\frac{z^{2}}{2}+\sigma(\eta)\,z+O(1).$
Since $\log\phi(z)=-\tfrac{1}{2}z^{2}-\tfrac{1}{2}\log(2\pi)$, we obtain
$\log\partial_{z}T_{h}(\eta,z)=\bigl[-\tfrac{1}{2}z^{2}+O(1)\bigr]\;-\;\bigl[\tfrac{1}{2}z^{2}+\sigma(\eta)\,z+O(1)\bigr]=-\,\sigma(\eta)\,z+O(1),$
uniformly in $\eta\in K_{\eta}$ as $|z|\to\infty$. With $\sigma_{\min}\leq\sigma(\eta)\leq\sigma_{\max}$ on $K_{\eta}$, this yields the two–sided bound
$\sigma_{\max}\,|z|-C\ \leq\ \log\partial_{z}T_{h}(\eta,z)\ \leq\ \sigma_{\max}\,|z|+C$
for some constant $C<\infty$ depending only on $K_{h}$, $K_{\eta}$, and the bounds on $m(\eta)$ and $\sigma(\eta)$. Equivalently,
$\bigl|\log\partial_{z}T_{h}(\eta,z)\bigr|\ \leq\ C_{0}+C_{1}\,|z|\qquad(|z|\ \text{large}),$
with $C_{1}$ proportional to $\sup_{\eta\in K_{\eta}}\sigma(\eta)$. This establishes the envelopes in Assumption 1(iii) and in turn implies $G(M)=O(e^{O}(M))$.
@@
Let $r:=|\theta-\mu|/\alpha$ and assume $\alpha(\eta)\in[\underline{\alpha},\overline{\alpha}]$, $\beta(\eta)\in[\underline{\beta},\overline{\beta}]$ with $\underline{\alpha}>0$, $2\geq\overline{\beta}$, $\underline{\beta}>0$, and $|\mu(\eta)|\leq M_{\mu}$ uniformly on $K_{\eta}$. The density is
@@
Hence, as $x\to\infty$,
$\bar{F}_{\eta}(x)\asymp\exp\!\Big(-\big(\tfrac{x-\mu}{\alpha}\big)^{\beta}\Big),$
with the precise leading factor $\tfrac{1}{2\,\Gamma(1/\beta)}\big(\tfrac{x-\mu}{\alpha}\big)^{1-\beta}$.
@@
At Gaussian tail level $\varepsilon_{N}=\Phi(-N)\sim\phi(M)/M$,
$\log\frac{1}{\varepsilon_{N}}=\tfrac{1}{2}M^{2}+\log M+\log\sqrt{2\pi},$
@@
$Q_{\eta}(1-\varepsilon_{N})\sim\alpha\Big(\tfrac{1}{2}M^{2}\Big)^{1/\beta},\qquad G(M)=\sup_{\eta\in K_{\eta}}Q_{\eta}(1-\varepsilon_{N})=O\!\big(M^{2/\underline{\beta}}\big).$
@@
Definition of $L_{S}(M)$. For $|\theta\,|\,\leq M$ and $\eta\in K_{\eta}$,
@@
Finally, as $\beta(\eta)$ is always less than or equals to two, the tail decay is always as thick as a Gaussian and part(iv) of the assumption is satisfied.
@@
$p(\theta\mid\eta)=\sum_{k=1}^{K}\pi_{k}(\eta)\,p_{k}(\theta\mid\eta),\qquad\pi_{k}(\eta)\geq 0,\ \sum_{k}\pi_{k}(\eta)=1,$
with component cdfs $F_{k}$, survival functions $\bar{F}_{k}$, and overall survival $\bar{F}=1-F$. For each component we define the envelope directly from the upper tail quantile as
$G_{k}(M)\;=\;\sup_{\eta\in K_{\eta}}\,\big|Q_{k,\eta}(1-\varepsilon_{N})\big|,\qquad\varepsilon_{N}=\Phi(-N)\sim\frac{\phi(M)}{M},$
where $Q_{k,\eta}(u)=F_{k,\eta}^{-1}(u)$ is the component quantile.
At the tail level, the mixture is governed by the component (or set of components) with the heaviest tail. Suppose there is a nonempty index set $H$ such that every component outside $H$ has vanishing relative tail mass compared to some $h\in H$, while components within $H$ have comparable tails. For large $x$, the survival of the mixture is
$\bar{F}(x\mid\eta)=\sum_{k=1}^{K}\pi_{k}(\eta)\,\bar{F}_{k}(x\mid\eta).$
Under the heaviest–tail assumption, the terms with $k\notin H$ vanish relative to those with $k\in H$, and the latter are comparable to each other. Hence
$\bar{F}(x\mid\eta)\ \sim\ \sum_{h\in H}\pi_{h}(\eta)\,\bar{F}_{h}(x\mid\eta)\ \sim\ C(\eta)\,\max_{k}\bar{F}_{k}(x\mid\eta),$
uniformly in $\eta$, where $C(\eta)$ lies between $\min_{h\in H}\pi_{h}(\eta)$ and $\sum_{h\in H}\pi_{h}(\eta)$.
To translate this into quantiles, note that for $\varepsilon>0$, the quantile $Q_{\eta}(1-\varepsilon)$ is defined by
$\bar{F}\!\big(Q_{\eta}(1-\varepsilon)\mid\eta\big)=\varepsilon.$
If $\bar{F}(x\mid\eta)\asymp\max_{k}\bar{F}_{k}(x\mid\eta)$, then the solution $x=Q_{\eta}(1-\varepsilon)$ must be of the same order as the largest among the component quantiles $Q_{k,\eta}(1-\varepsilon)$. Thus, at Gaussian tail levels $\varepsilon_{N}=\Phi(-N)$,
$Q_{\eta}(1-\varepsilon_{N})\ \asymp\ \max_{k}Q_{k,\eta}(1-\varepsilon_{N}),\qquad G(M)\ \asymp\ \max_{k}G_{k}(M).$
@@
$L_{S}(M)\ \leq\ \max_{k}M_{\theta,k}(M),\qquad L_{CP}(M)\ \leq\ \max_{k}M_{\eta,k}(M)+\sup_{k}\|\nabla_{\eta}\log\pi_{k}\|_{L^{\infty}(K_{\eta})}.$
Assume the mixing weights $\pi_{k}:K_{\eta}\to(0,1)$ are continuously differentiable and bounded below by $\pi_{\min}>0$. Then

### v2

@@ Acknowledgements
$\begin{split}L_{S}(M)=\sup_{\eta\in K_{\eta}}\ \sup_{\begin{subarray}{c}\theta\neq\theta^{\prime}\\ |\theta\,|\,,|\theta^{\prime}|\leq M\end{subarray}}\frac{|\log p(\theta\mid\eta)-\log p(\theta^{\prime}\mid\eta)|}{|\theta-\theta^{\prime}|},\\ L_{CP}(M)=\sup_{|\theta\,|\,\leq M}\ \sup_{\eta\neq\eta^{\prime}}\frac{|\log p(\theta\mid\eta)-\log p(\theta\mid\eta^{\prime})|}{|\eta-\eta^{\prime}|_{1}}.\end{split}$
For $G(M)$ we evaluate upper envelopes of $Q_{\eta}$ at the Gaussian tail levels $\varepsilon_{M}:=\bar{\Phi}(M)=1-\Phi(M)\sim\phi(M)/M$. On $K_{\eta}$ (compact), continuous parameter maps are bounded; for $L_{CP}(M)$ we use that $\log p(\theta\mid\eta)$ is smooth in its parameters on compact sets and apply the chain rule with the (finite) Lipschitz moduli of the parameter maps induced by continuity on $K_{\eta}$. [^2] For standard parametric families, the derivatives of $\log p$ with respect to the parameters are bounded on compacts; combined with bounded variation of the parameter maps on $K_{\eta}$, this gives the displayed $L_{CP}(M)$ orders.
@@
$\begin{split}Q_{\eta}(u)&=\mu(\eta)+\sigma(\eta)\Phi^{-1}(u)\ \Rightarrow\ T(\eta,z)=\mu(\eta)+\sigma(\eta)z,\\ G(M)&=\sup_{\eta\in K_{\eta}}\,|\,\mu(\eta)|+\left(\sup_{\eta\in K_{\eta}}\sigma(\eta)\right)\,M=O(M).\end{split}$
$p\left(\theta\,|\,\mu,\sigma\right)=\frac{1}{\sigma\sqrt{2\pi}}\exp\left[-\frac{(\theta-\mu)^{2}}{2\sigma^{2}}\right].$
$\log p\left(\theta\,|\,\mu,\sigma\right)=\text{constant}-\log\sigma-\frac{(\theta-\mu)^{2}}{2\sigma^{2}}.$
$\partial_{\theta}\log p=-\frac{(\theta-\mu)}{\sigma^{2}}\ \Rightarrow\ L_{S}(M)=O\left(\frac{M}{\sigma^{2}_{\min}}\right)=O(M).$
@@
For $Z\sim\mathcal{N}(0,1)$, by Mill’s ratio, the upper Gaussian tail satisfies $\varepsilon_{M}=\bar{\Phi}(M)=1-\Phi(M)\sim\phi(M)/M$.
@@
So, $Q_{\eta}(\Phi(z))=O(1/\bar{\Phi}(z))=z\exp(z^{2}/2)$ as $z\to\infty$, giving the bound
$G(M)=O\!\left(Me^{M^{2}/2}\right).$
$\log p(\theta\mid\mu,\gamma)=\text{constant}+\log\gamma-\log\left(\gamma^{2}+(\theta-\mu)^{2}\right).$
The log-density derivatives are
$\partial_{\theta}\log p=-\frac{2(\theta-\mu)}{\gamma^{2}+(\theta-\mu)^{2}},\qquad\partial_{\mu}\log p=\frac{2(\theta-\mu)}{\gamma^{2}+(\theta-\mu)^{2}},\qquad\partial_{\gamma}\log p=\frac{1}{\gamma}-\frac{2\gamma}{\gamma^{2}+(\theta-\mu)^{2}}.$
Using $2ab\leq a^{2}+b^{2}$, these satisfy the uniform bounds
@@
Student-$t_{\nu(\eta)}$ with location $\mu(\eta)$, scale $s(\eta)$.
Let $T\sim t_{\nu}$ with density
$f_{\nu}(u)=c_{\nu}\Bigl(1+\tfrac{u^{2}}{\nu}\Bigr)^{-(\nu+1)/2},\qquad c_{\nu}=\frac{\Gamma\!\bigl(\frac{\nu+1}{2}\bigr)}{\sqrt{\nu\pi}\,\Gamma\!\bigl(\frac{\nu}{2}\bigr)}.$
For $x>0$,
$\overline{F}_{\nu}(x)=\mathbb{P}(T>x)=\int_{x}^{\infty}f_{\nu}(u)\,du=c_{\nu}\int_{x}^{\infty}\Bigl(1+\tfrac{u^{2}}{\nu}\Bigr)^{-(\nu+1)/2}\,du.$
As $u\to\infty$,
$\Bigl(1+\tfrac{u^{2}}{\nu}\Bigr)^{-(\nu+1)/2}=\Bigl(\tfrac{u^{2}}{\nu}\Bigr)^{-(\nu+1)/2}\!\Bigl(1+O(u^{-2})\Bigr)=\nu^{(\nu+1)/2}u^{-(\nu+1)}\Bigl(1+O(u^{-2})\Bigr),$
hence
$\overline{F}_{\nu}(x)=c_{\nu}\nu^{(\nu+1)/2}\!\int_{x}^{\infty}u^{-(\nu+1)}\Bigl(1+O(u^{-2})\Bigr)\,du=c_{\nu}\nu^{(\nu+1)/2}\!\left(\frac{x^{-\nu}}{\nu}+O\!\left(x^{-(\nu+2)}\right)\right).$
With
$C_{\nu}\coloneqq c_{\nu}\nu^{(\nu-1)/2}=\frac{\Gamma\!\bigl(\frac{\nu+1}{2}\bigr)}{\sqrt{\pi}\,\Gamma\!\bigl(\frac{\nu}{2}\bigr)}\,\nu^{(\nu-1)/2},$
this gives
$\overline{F}_{\nu}(x)=C_{\nu}x^{-\nu}\Bigl(1+O(x^{-2})\Bigr),\qquad x\to\infty.$
For a location scale family $X=\mu+sT$ with fixed $\mu\in\mathbb{R}$ and $s>0$,
$\overline{F}_{\mu,s}(x)=\mathbb{P}(X>x)=\overline{F}_{\nu}\!\left(\frac{x-\mu}{s}\right)\sim C_{\nu}s^{\nu}(x-\mu)^{-\nu}\sim c_{\nu}x^{-\nu},\qquad x\to\infty,$
where $c_{\nu}=C_{\nu}s^{\nu}$. Tail $\bar{F}_{\eta}(x)\sim c_{\nu(\eta)}\,x^{-\nu(\eta)}$ yields
$Q_{\eta}(1-\varepsilon)\sim(c_{\nu(\eta)}/\varepsilon)^{1/\nu(\eta)}\ \Rightarrow\ G(M)=O\left(\left(\frac{M}{\phi(M)}\right)^{1/\nu_{\min}}\right)=O\left(\left(Me^{M^{2}/2}\right)^{1/\nu_{\min}}\right),$
@@
$p(\theta\mid\nu,\mu,s)=\frac{\Gamma\left(\frac{\nu+1}{2}\right)}{s\sqrt{\pi\nu}\,\Gamma\left(\frac{\nu}{2}\right)}\,\left(1+\frac{(\theta-\mu)^{2}}{\nu s^{2}}\right)^{-\frac{\nu+1}{2}}.$
$\log p(\theta\mid\nu,\mu,s)=\text{constant}-\log s+\log\Gamma\left(\frac{\nu+1}{2}\right)-\log\Gamma\left(\frac{\nu}{2}\right)-\frac{1}{2}\log\nu-\frac{\nu+1}{2}\log\left(1+\frac{(\theta-\mu)^{2}}{\nu s^{2}}\right).$
$\partial_{\theta}\log p=-\frac{(\nu+1)(\theta-\mu)}{\nu s^{2}+(\theta-\mu)^{2}}=-\left(\frac{\nu+1}{2\sqrt{\nu}s}\right)\cdot\frac{2\sqrt{\nu}s\,(\theta-\mu)}{\nu s^{2}+(\theta-\mu)^{2}}\ \Rightarrow\ L_{S}(M)=O(1).$
Derivatives w.r.t. $(\mu,s)$ are bounded following similar arguments. The $\nu$-term contains $\tfrac{1}{2}\log\!\big(1+(\theta-\mu)^{2}/(\nu s^{2})\big)$, hence on [M,M][-M,M] gives $O(\log M)$:
@@
With normal base $Z\sim\mathcal{N}(0,1)$, let $\varepsilon_{M}=\bar{\Phi}(Z)\sim\phi(M)/M$. Then
@@
Log-density derivatives.
$\log p\left(\theta\mid\eta\right)=\text{constant}-\log b-\frac{|\theta-\mu|}{b}.$
For $\theta\neq\mu$,
@@
Uniform bounds on the $\theta$-truncated set $\{|\theta-\mu|\leq M\}$:
@@
Symmetrized mollified log-normal.
@@
Let $\mu(\eta),\sigma(\eta)$ be continuous on compact $K_{\eta}$ with $0<\sigma_{\min}\leq\sigma(\eta)\leq\sigma_{\max}<\infty$. Define the symmetrized log-normal density
$f_{\mathrm{SLN}}(\theta\mid\eta)=\tfrac{1}{2}f_{\mathrm{LN}}\left(\left|\theta\right|\mid\eta\right),\qquad f_{\mathrm{LN}}(x\mid\eta)=\frac{1}{x\sigma(\eta)\sqrt{2\pi}}\exp\left(-\frac{\left(\log x-\mu(\eta)\right)^{2}}{2\sigma^{2}(\eta)}\right),\quad x>0.$
Let $K_{h}$ be a symmetric $C^{\infty}$ mollifier with compact support [h,h][-h,h], $\int K_{h}=1$, and $K_{h}(u)\geq\kappa_{0}>0$ for $|u|\leq h/2$. Define the smoothed symmetric log-normal
@@
Since $f_{\mathrm{LN}}(\cdot\mid\eta)$ is decreasing for $x$ beyond its mode, the worst case over $|\theta\,|\,\leq M$ is attained at $|\theta\,|\,=M$. Thus, for all $\eta\in K_{\eta}$,
$\inf_{|\theta\,|\,\leq M}p_{h}(\theta\mid\eta)\ \geq\ \frac{\kappa_{0}}{2}\!\int_{M-h/2}^{M+h/2}f_{\mathrm{LN}}(x\mid\eta)\,dx\ \geq\ \frac{\kappa_{0}h}{4}\,\inf_{\eta\in K_{\eta}}\ \inf_{x\in[M-h/2,\,M+h/2]}f_{\mathrm{LN}}(x\mid\eta).$
@@
$L_{S}(M)$ $=O\Big((M+1)\exp\Big(\tfrac{(\log(M+h/2)-\mu_{\max})^{2}}{2\sigma_{\min}^{2}}\Big)\Big)=O(Me^{O\left((\log M)^{2}\right)}),$
$L_{CP}(M)$ $=O\Big((\log M)^{2}(M+1)\exp\Big(\tfrac{(\log(M+h/2)-\mu_{\max})^{2}}{2\sigma_{\min}^{2}}\Big)\Big)=O((\log M)^{2}Me^{O\left((\log M)^{2}\right)}).$
Quantile growth Let $X\sim f_{\mathrm{SLN}}(\cdot\mid\eta)$ and $Y\sim K_{h}$ with $\operatorname{supp}(Y)\subset[-h,h]$, independent. Let $F_{X},F_{X+Y}$ be the distribution functions and $Q_{X},Q_{X+Y}$ their generalized quantile functions $Q(u):=\inf\{x:\,F(x)\geq u\}$, $u\in(0,1)$.
For every $x\in\mathbb{R}$,
$F_{X+Y}(x)=\mathbb{E}\!\left[F_{X}(x-Y)\right]\in\big[\,F_{X}(x-h),\,F_{X}(x+h)\,\big],$
because $F_{X}$ is nondecreasing and $Y\in[-h,h]$ almost surely.
By monotonicity of $F_{X}$ and of its generalized inverse, for all $u\in(0,1)$,
$Q_{X}(u)-h\;\leq\;Q_{X+Y}(u)\;\leq\;Q_{X}(u)+h,\qquad\text{hence}\qquad\big|Q_{X+Y}(u)-Q_{X}(u)\big|\leq h.$
If $X\sim\mathrm{LN}(m(\eta),\sigma(\eta))$ with $m,\sigma$ bounded on $K_{\eta}$ and $\sigma(\eta)\geq\sigma_{\min}>0$, then
$Q_{X}\!\left(\Phi(z)\right)=\exp\!\big(m(\eta)+\sigma(\eta)z\big).$
Combining with the previous display yields
$Q_{X+Y}\!\left(\Phi(z)\right)=\exp\!\big(m(\eta)+\sigma(\eta)z\big)\,e^{O(1)}\quad\text{uniformly in }\eta\in K_{\eta}.$
Therefore, for $|z|\leq M$,
$Q_{X+Y}\!\left(\Phi(z)\right)=\exp\!\big(O(M)\big)\quad\text{uniformly in }\eta\in K_{\eta},$
with constants depending only on $h$ and the bounds for $m$ and $\sigma$.
@@
Let $r:=|\theta-\mu|/\alpha$. By continuity, we have $\alpha(\eta)\in[\underline{\alpha},\overline{\alpha}]$, $\beta(\eta)\in[\underline{\beta},\overline{\beta}]$ with $\underline{\alpha}>0$, $\underline{\beta}>0$, $2\geq\overline{\beta}$, and $|\mu(\eta)|\leq M_{\mu}$ uniformly on $K_{\eta}$. The density is
@@
Let
$y:=\left(\frac{x-\mu}{\alpha}\right)^{\beta},\qquad C:=\frac{1}{2\,\Gamma(1/\beta)}.$
Then
$\bar{F}_{\eta}(x)\sim C\,y^{\frac{1-\beta}{\beta}}e^{-y}.$
Solving $\bar{F}_{\eta}(x)\sim\varepsilon$ for $\varepsilon\downarrow 0$ gives
$y=\log\frac{1}{\varepsilon}+\frac{1-\beta}{\beta}\log\!\Big(\log\frac{1}{\varepsilon}\Big)+\log C+o(1).$
Hence the $(1-\varepsilon)$-quantile satisfies
$Q_{\eta}(1-\varepsilon)=\mu+\alpha\left[\log\frac{1}{\varepsilon}+\frac{1-\beta}{\beta}\log\!\Big(\log\frac{1}{\varepsilon}\Big)+\log\!\Big(\frac{1}{2\,\Gamma(1/\beta)}\Big)+o(1)\right]^{1/\beta}.$
@@
At Gaussian tail level $\varepsilon_{M}=\Phi(-Z)\sim\phi(M)/M$,
$\log\left(\frac{1}{\varepsilon_{M}}\right)=\tfrac{1}{2}M^{2}+\log M+\log\sqrt{2\pi},$
@@
$Q_{\eta}(1-\varepsilon_{N})\sim\alpha\Big(\tfrac{1}{2}M^{2}\Big)^{1/\beta},\qquad G(M)=\sup_{\eta\in K_{\eta}}Q_{\eta}(1-\varepsilon_{M})=O\big(M^{2/\underline{\beta}}\big).$
@@
Here $\psi$ is the digamma function. Definition of $L_{S}(M)$. For $|\theta\,|\,\leq M$ and $\eta\in K_{\eta}$,
@@
Finally, as $\beta(\eta)$ is always less than or equal to two, the tail decay is always as thick as a Gaussian and part(iv) of the assumption is satisfied.
@@
$p(\theta\mid\eta)=\sum_{k=1}^{K}\pi_{k}(\eta)\,p_{k}(\theta\mid\eta),\qquad\pi_{k}(\eta)\geq 0,\ \sum_{k}\pi_{k}(\eta)=1.$
[…]
@@
$L_{S}(M)\ \leq\ \max_{k}L_{S,k}(M),\qquad L_{CP}(M)\ \leq\ \max_{k}L_{CP,k}(M)+\sup_{k}\|\nabla_{\eta}\log\pi_{k}\|_{L^{\infty}(K_{\eta})}.$
Assume the mixing weights $\pi_{k}:K_{\eta}\mapsto(0,1)$ are continuously differentiable and bounded below by $\pi_{\min}>0$. Then